Fermi National Laboratory
 

Accelerator Update for June 1, 2007 - June 4, 2007

 

Friday June 1

The day shift began with Operations monitoring Tevatron (TeV) store 5469 (113.33E30) and the Recycler stash (114.1E10), with the Antiproton Source (Pbar) stacking (45.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 9:24 AM, a Linac Klystron RF station (KRF5) tripped off. It reset.

A TeV expert set up the #1 Tevatron Electron Lens (TEL1) at 9:32 AM to temporarily clean the abort gap so that he could take TEL2 down for repairs. After the repair both devices were set back to their normal operation.

A MI expert began a slipstacking study at 11:21 AM.

Pbar experts began a reverse proton study at 12:40 PM.

Operations shut down Linac at 1:03 PM for LRF3 repairs and an access into the Preaccelerator dome. Booster experts began work on a power supply (GMPS).

Linac experts completed their work on LRF3 at 2:30 PM. Pbar resumed stacking.

At 3:14 PM, two ODH heads in Switchyard (SY) enclosure C began alarming. The alarm was caused by high humidity and not an oxygen deficiency. A tunnel ventilation fan had failed. The Duty Mechanic began repairs on the fan.

Operations reported at 4:22 PM that the Recycler had stashed antiprotons two times during the day shift.

At 4:52 PM, the Duty Mechanic reported that they didn’t have the parts to complete SY enclosure C the fan repair. A Safety Officer restricted access to the enclosure until the ventilation gets repaired.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday June 2

The midnight shift began with Operations monitoring store 5469 (47.06E30) and the stash (294.7E10), with Pbar stacking (3.4E10), and with NuMI and MiniBooNE taking beam.

At 3:39 AM, KRF5 tripped off. It reset.

At 6:04 AM, a NuMI kicker (I:KPS6N) tripped off and took down the NuMI beam permit. It reset.

At 6:49 AM, Operations reported that the MI beam was aborting due to a monitor that checked for bad beam positions, but the value the monitor was reporting was bad. Operations rebooted its VME and resolved the problem.

Operations terminated store 5469 at 8:01 AM.

Operations reported at 8:02 AM that the Recycler had stashed three times during the midnight shift.

Operations put the TeV into shot setup at 8:14 AM.

At 10:19 AM, KRF6 tripped off on a closed LCW valve.

Operations established store 5470 at 10:32 AM with an initial luminosity of 246.17E30.

The temperature regulation for KRF6 returned to normal and the valve opened at 10:59 AM.

Linac resumed sending beam to MiniBooNE, NuMI, and Pbar at 11 AM.

Operations reported at 8:01 PM that the Recycler had stashed five times during the last 12 hours. Accelerator status: store 5470 was at 91.92E30, the stack was at 37.0E10, and the stash was at 163.8E10.

Sunday June 3

The midnight shift began with Operations monitoring store 5470 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 6:57 AM, TEL2 tripped off. It reset.

Operations reported at 7:52 AM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5470 was at 50.84E30 and the stash was at 304.9E10.

Pbar experts began an Accumulator orbit adjustment and a Pbar target scan at 8:17 AM.

Operations terminated store 5470 at 12:01 PM.

Operations put the TeV into shot setup at 12:13 PM.

Operations established store 5471 at 2:09 PM with an initial luminosity of 257.18E30.

At 6:22 PM, KRF5 tripped off for the first of two times in about 35 minutes. It reset.

At 7:09 PM, Operations received a report of an LCW leak in Booster. Operators responded and found LCW leaking from the BRF19 station. Operators bypassed BRF19 and turned on BRF7 to make up for BRF19 being off.

At 7:31 PM, KRF2 tripped off for the first of multiple times over the next 45 minutes. Operators kept resetting the station.

Operations reported at 7:57 PM that the Recycler had stashed five times during the last 12 hours. Accelerator status: store 5471 was at 118.95E30.

At 8:21 PM, KRF2 tripped off. Operations contacted experts and then made adjustments to its switchtube balance.

At 9:52 PM, KRF2 tripped off. It reset.

Monday June 4

The midnight shift began with Operations monitoring store 5471 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 3:56 AM, BRF18 tripped off. It reset.

At 4:16 AM, BRF18 tripped and would not reset due to sparking. Operators bypassed the station. Operations lowered the Booster’s beam intensity.

At 4:47 AM, LRF2 tripped off and would not reset. Operators investigated and found its driver was off.

Operators reported at 5:26 AM that they had manually adjusted the filament timers on LRF2 and got the station on.

At 6:42 AM, a Booster vacuum pump (B:IPRF17) tripped off and brought down BRF7 & 8. Everything reset.

At 7:05 AM, a Booster vacuum pump (B:IPRF24) tripped off and brought down BRF15 & 16. Everything reset.

A Booster RF expert reported that he brought BRF18 online at 7:14 AM. Operations brought the Booster beam intensity back up to normal.

Operations reported at 8:09 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to terminate store 5471 around 10 AM and then allow CDF and D0 a three-hour access. Operations will put the TeV into shot setup after the access.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/04/2007    email Fermilab